# Optimal Attention Temperature Enhances In-Context Learning under Distribution Shift

### v1

@@ 4.4 Optimal attention temperature improves performance
Suppose Assumptions 3.1, 3.2, and 4.5 hold. To minimize the generalization error, the optimal attention temperature for inference is given by
$\tau_{\text{optimal}}=\frac{2\text{Tr}\left({\bm{A}}{\bm{M}}_{11}^{T}{\bm{F}}_{1}{\bm{M}}_{11}\right)}{\text{Tr}\left({\bm{A}}\left({\bm{F}}_{2}{\bm{M}}_{11}+{\bm{M}}_{11}^{T}{\bm{F}}_{2}^{T}\right)\right)},$ (15)
@@
We minimize the generalization error from Theorem 4.6 with respect to $\tau$ (Appendix I).
Consider the optimal temperature $\tau_{\text{optimal}}$ from Theorem 4.7. When $\tau_{\text{optimal}}\neq 1$, using an unadjusted temperature leads to suboptimal generalization error. Thus, incorporating the optimal temperature improves generalization in in-context learning under distribution shift.
A natural question is whether the optimal temperature can completely mitigate the adverse effects of distribution shifts. This depends on both the pretraining and test distributions. In some settings, the adjustment can entirely compensate for the shift. For example, if the task distribution is fixed as ${\bm{w}}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$, the noise variance is $\sigma=0$, and the input distribution changes from ${\bm{x}}_{\text{train}}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$ to ${\bm{x}}_{\text{test}}\sim\mathcal{N}(\mathbf{0},c{\bm{I}})$, then the optimal temperature $\tau_{\text{optimal}}=c$ fully counteracts the shift. In more complex scenarios, it may only partially mitigate the impact, yet still yields improved ICL.
@@
5 Experimental results
In this section, we empirically validate our theory and show that optimal attention temperature consistently enhances generalization. We begin with controlled linear regression experiments using (i) the simplified Transformer model with linearized attention (5) and (ii) GPT-2 (Radford et al., 2019), which combines multi-head softmax attention with MLP layers [^2] GPT-2 results are in Appendix K.. These experiments confirm that our theoretical insights transfer from simplified to expressive architectures. Finally, we evaluate Llama2-7B (Touvron et al., 2023) on SCIQ in-context learning tasks (Welbl et al., 2017), demonstrating that temperature selection is a principled and effective lever for improving robustness in large language models.
5.1 Experiments on linear regression tasks
We consider a Transformer architecture with linearized attention and no MLP layers, as analyzed in our theoretical development. Figures 1 and 2 illustrate its behavior on linear regression tasks (2). Theoretical predictions closely match empirical performance across a range of conditions, confirming the robustness of our analysis. In Figure 1, we compare the ICL performance of the model with and without applying the optimal temperature. As context length $l$ increases (Figure 1a), the model’s predictions converge to those of the Bayes-optimal linear model, validating its ICL capability. Figure 1b shows that under an input covariance shift, model performance degrades—but applying the optimal temperature restores alignment with the Bayes-optimal solution. Additionally, Figure 1c shows that the influence of task distribution shift decreases as $l$ increases.
@@ 5.1 Experiments on linear regression tasks
Figure 3: Effect of attention temperature on the ICL performance of LLaMA-2-7B (Touvron et al., 2023) on the SCIQ dataset (Welbl et al., 2017). Distribution shift is induced by injecting noisy yet “relevant” labels into in-context demonstrations following Gao et al. (2024). Panel (a) fixes the noisy ratio at 0.6; panel (b) fixes the number of in-context examples at 6. Results (averaged over 12 Monte Carlo runs) include error bars showing one standard deviation. Attention temperature of all the layers is set to $\tau\sqrt{d_{k}}$ for dimension independence, where $d_{k}$ denotes the key dimension of the corresponding layer. Furthermore, the dashed black line marks the “optimal temperature” computed from the variance-to-mean ratio of pre-softmax scores, which is an insight derived from Theorem 4.7, as explained in Appendix J. Full experimental details appear in Appendix K.
5.2 Experiments with LLMs for in-context question answering tasks
@@
Figure 3 shows the results. In (a), performance versus the number of in-context examples under fixed noise reveals a non-monotonic trade-off between added context and accumulated noise. In (b), higher noise ratios push the optimal temperature upward, matching our theoretical prediction (cf. Figure 2c). Together, these experiments demonstrate that the optimal temperature is not only theoretically motivated but also an effective tool for improving ICL robustness in real-world LLMs.
@@ 6 Conclusion
This work provides a unified theoretical and empirical account of how attention temperature governs the in-context learning (ICL) performance of pretrained Transformers under distribution shift. Using a simplified yet expressive framework based on linearized softmax attention, we analytically show how shifts in input covariance and label noise degrade ICL and derive an optimal temperature that provably minimizes generalization error. Extensive experiments on synthetic regression tasks, GPT-2, and LLaMA-2 validate our predictions, demonstrating that temperature selection is not a mere heuristic but a principled mechanism for improving robustness. Taken together, our results advance the theoretical understanding of Transformer behavior under distribution shift and establish attention temperature as a powerful, practical lever for building more adaptive and generalizable foundation models.
Acknowledgments
We acknowledge that this work is supported partially by TÜBİTAK under project 124E063 in ARDEB 1001 program. We extend our gratitude to TÜBİTAK for its support. S.D. is supported by an AI Fellowship provided by KUIS AI Center and a PhD Scholarship (BİDEB 2211) from TÜBİTAK.
Ethics statement
Our research, as presented in the paper, conforms with the Code of Ethics. As the work is mostly of a theoretical nature, it does not involve any component that can be subject to ethical concerns.
Reproducibility statement
All results and settings are explained in detail to streamline reproducibility. Derivations of the theoretical results can be found in the appendix. The settings used to generate the figures are explained in Section 3, in the corresponding captions, and in the appendix. Finally, the code for the experimental results will be released with the camera-ready version of this work.
The use of Large Language Models (LLMs)
During the writing of the paper, we utilize LLMs in order to sharpen the presentation language. In doing so, we provide a sentence or a paragraph that we wrote before, instruct an LLM model to rewrite it in a better tone, and use the produced sentence/paragraph whenever it clearly describes what we aim to describe. This approach has been repeatedly applied to multiple parts of the paper to refine the writing. Overall, we take full responsibility for the contents written.
Appendix A Derivation of Bayes-optimal ridge estimator for $\bm{w}$}
@@
Appendix B Derivation of linearized softmax
@@
To obtain a linear approximation, we expand around the origin ${\bm{z}}=\mathbf{0}$ using a first-order Taylor series:
@@
Substituting back, we obtain the linearized softmax:
@@
$=:\text{linearized\_softmax}({\bm{z}}).$ (34)
This derivation yields the linearized attention formulation in Eq. (5). From a practical standpoint, linearized attention mechanisms have been empirically evaluated and shown to achieve performance comparable to standard softmax attention (Han et al., 2024).
Appendix C Linear vs. linearized attention for in-context learning
Here, we highlight the distinction between linear attention and linearized attention in the context of the linear regression problem defined in Eq. (2). Analytically, the key difference lies in the fact that the linearized attention model operates on centered input data, whereas the linear attention model uses raw data without centering. Apart from this centering step, both mechanisms are equivalent, except that linearized attention includes an additional bias term ($b_{Att}$ in Eq. (10)). However, this bias term is inconsequential in our linear regression setting and does not affect the predictive outcome.
Thus, the data-centering operation is the principal differentiator in our analysis. Specifically, linear attention’s omission of centering makes it sensitive to shifts in the input mean, whereas linearized attention remains robust under such transformations. We illustrate this phenomenon in Figure 4, where we simulate a shift in the input mean at test time. The results demonstrate that linear attention fails to recover Bayes-optimal performance under mean shift, indicating its limitations for in-context learning in this setting. In contrast, linearized attention successfully compensates for the mean shift and achieves Bayes-optimal performance as the number of context points $l$ increases.
Therefore, in the presence of possible distributional shifts—particularly in the input mean—linearized attention offers a more robust and theoretically grounded alternative to linear attention for in-context learning.
Figure 4: Comparison of linear and linearized attention under a shift in input mean. The plot illustrates the impact of a test-time shift in input mean on the performance of linear attention and linearized attention. While linear attention degrades under the distribution shift and fails to recover the Bayes-optimal performance, linearized attention remains robust and asymptotically matches the Bayes-optimal predictor as the number of context length $l$ increases.
Appendix D Temperature effects for softmax and linearized softmax
@@
In the linearized case, temperature similarly acts as an inverse scaling of the variance of the output components, capturing the limit $\tau\to\infty$ (all elements equal to $1/l$). For $\tau\to 0^{+}$, linearized softmax also reflects the maximal variance, but it does not produce a true one-hot distribution. Thus, linearized softmax closely mirrors the temperature behavior of softmax, except in the degenerate limit $\tau\to 0^{+}$, which is not of practical relevance in this work.
To further illustrate these effects, Figure 5 compares softmax and linearized softmax across different temperatures. The figure demonstrates that linearized softmax faithfully captures the variance effect of temperature: the variance of the output components is inversely proportional to $\tau$. Moreover, as $\tau\to\infty$, both softmax and linearized softmax concentrate around $1/l$, whereas linear attention with temperature scaling does not. Overall, the output distributions of softmax and linearized softmax are highly similar, except at very small values of $\tau$, where linearized softmax may yield negative components while softmax tends toward sparsity with many zeros. By contrast, linear attention with temperature scaling produces qualitatively different distributions. This comparison highlights the advantage of linearized softmax as a faithful surrogate for analyzing temperature effects relevant to softmax.
Figure 5: Comparison of temperature effects of softmax, linearized softmax, and linear (with temperature scaling) cases. We consider an input vector ${\bm{x}}\in\mathbb{R}^{l}$ whose histogram is illustrated on the left-most plot. Rest of the plots illustrates histograms of the elements of $\text{softmax}({\bm{x}}/\tau)$, $\text{linearized\_softmax}({\bm{x}}/\tau)$ defined in (34) and ${\bm{x}}/(l\tau)$ from left to right, respectively.
Appendix E Expanded form of linearized attention
Using block matrix notation, the prediction from the linearized attention model can be expanded as:

### v2

@@ 4.2 Optimal attention temperature
Suppose Assumptions 3.1, 3.2, and 4.1 hold. To minimize the generalization error, the optimal attention temperature for inference is given by
$\tau_{\text{opt}}=\frac{2\text{Tr}\left({\bm{A}}{\bm{M}}_{11}^{T}{\bm{F}}_{1}{\bm{M}}_{11}\right)}{\text{Tr}\left({\bm{A}}\left({\bm{F}}_{2}{\bm{M}}_{11}+{\bm{M}}_{11}^{T}{\bm{F}}_{2}^{T}\right)\right)},$ (16)
@@
We minimize the generalization error from Theorem 4.2 with respect to $\tau$ (Appendix I). ∎
Consider the optimal temperature $\tau_{\text{opt}}$ from Theorem 4.3. When $\tau_{\text{opt}}\neq 1$, using an unadjusted temperature leads to suboptimal generalization error. Thus, incorporating the optimal temperature improves generalization in in-context learning under distribution shift.
@@
4.3 Effect of distribution shift
Pretraining for explaining distribution shift effects —
To streamline our explanations below, for pretraining, we optimize the parameters $\bm{V}$} and $\bm{M}$} using $m$ samples of $({\bm{Z}},y_{l})$ drawn from the distribution $\mathcal{D}^{train}$, where each $\bm{Z}$} contains l1l-1 $({\bm{x}},y)$ pairs intended for ICL. Building upon prior work that connects ICL in linear regression to the Bayes-optimal ridge estimator (Zhang et al., 2024; Raventós et al., 2023), we configure $\bm{M}$} and $\bm{V}$} to emulate Bayes-optimal ridge regression. Specifically, we aim for $\hat{{\bm{w}}}_{Att}({\bm{C}}_{xx},{\bm{C}}_{xy};{\bm{M}},{\bm{V}})\approx\hat{{\bm{w}}}_{Bayes}$ and $b_{Att}({\bm{s}}_{x},s_{y};{\bm{V}})\approx 0$.
Proposition 4.4 (Pretrained parameters).
Remark 4.5.
While the pretrained parameters specified in Proposition 4.4 are not guaranteed to be optimal in all settings, they are analytically useful for examining the effects of distribution shifts.
Based on Proposition 4.4, we arrive at the following:
Corollary 4.6.
Suppose there is no distribution shift between training and inference. Then, under the parameter configuration of Proposition 4.4, the Transformer model (7) emulates the Bayes-optimal linear model, implying that it is capable of in-context learning according to Definition $3.$
ICL under distribution shift —
Case I: Shift in input distribution —
  1. (i)
    Mean shift (${\bm{\mu}}_{x}^{train}\neq{\bm{\mu}}_{x}^{test}$): Centering renders the approximate model invariant to mean shifts, but the uncentered linear attention model remains sensitive, as noted in Remark 3.4.
  2. (ii)
Case II: Shift in task distribution —
The task vectors follow ${\bm{w}}\sim\mathcal{N}({\bm{\mu}}_{w},{\bm{\Sigma}}_{w})$. Let ${\bm{\mu}}_{w}^{train},{\bm{\Sigma}}_{w}^{train}$ and ${\bm{\mu}}_{w}^{test},{\bm{\Sigma}}_{w}^{test}$ be the mean and covariance of the task distribution during pretraining and testing, respectively. The Transformer model can incorporate ${\bm{\mu}}_{w}^{train}$ and ${\bm{\Sigma}}_{w}^{train}$ via the pretrained parameters ${\bm{M}}_{11}$ and ${\bm{v}}_{21}$ (see Proposition 4.4). However, as the context length $l$ increases, the model’s dependence on the task distribution diminishes. Thus, shifts in the task distribution primarily affect ICL performance for small $l$.
Case III: Shift in noise distribution —
Summary — The Transformer model is robust to shifts in input mean but sensitive to input covariance changes. Shifts in task or noise distribution reduce ICL performance at small $l$, though increasing $l$ mitigates these effects.
4.4 Interpretation of the optimal attention temperature (Theorem 4.3)
To obtain a simplified expression, we consider a simple but representative family of shifts as follows. The training distributions are
${\bm{x}}_{i}\sim\mathcal{N}(\mathbf{0},{\bm{I}}),\qquad{\bm{w}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}),\qquad\epsilon_{i}\sim\mathcal{N}(0,\hat{\sigma}^{2}).$
We then introduce three independent shift parameters for the test distribution:
${\bm{x}}_{i}\sim\mathcal{N}(\mathbf{0},a{\bm{I}}),\qquad{\bm{w}}\sim\mathcal{N}(\mathbf{0},b{\bm{I}}),\qquad\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}),$
where $a>0$ controls the input variance shift, $b>0$ controls the task-parameter variance shift, and $\sigma>0$ controls the noise-variance shift. This setting preserves isotropy, which makes it possible to derive a clean closed-form expression while still connecting directly to realistic distribution shifts.
Substituting these shifted distributions into the optimal-temperature expression in (16) yields
$\tau_{\text{opt}}$ $=\frac{2\text{Tr}\left(a{\bm{I}}{\bm{M}}_{11}^{T}\left(ab{\bm{I}}+\frac{1}{l}(\sigma^{2}+abd){\bm{I}}\right)a{\bm{I}}{\bm{M}}_{11}\right)}{\text{Tr}\left(a{\bm{I}}\left(ab{\bm{I}}\left({\bm{M}}_{11}+{\bm{M}}_{11}^{T}\right)\right)\right)},$ (18)
$=\left(a+\frac{1}{l}\left(\frac{\sigma^{2}}{b}+ad\right)\right)\frac{\text{Tr}\left({\bm{M}}_{11}^{T}{\bm{M}}_{11}\right)}{\text{Tr}\left({\bm{M}}_{11}\right)}.$ (19)
This concrete formula makes several effects fully explicit:
  • Input shift — Increasing input variance $a$ directly scales $\tau_{\text{opt}}$ upward. This aligns with our earlier results (Figure 2) and the heuristic derived in Appendix J, which suggests that a greater variance of pre-softmax scores requires a higher temperature to maintain robustness to input shifts.
  • Noise shift — Increasing noise variance $\sigma^{2}$ also increases $\tau_{\text{opt}}$, but only through the $\tfrac{1}{l}$ term, reflecting the diminishing effect of noise when more in-context examples are available.
  • Task shift — Increasing task variance $b$ reduces the effect of noise (via $\sigma^{2}/b$), slightly lowering the optimal temperature.
  • Context length — As $l\to\infty$, the $\tfrac{1}{l}$ term vanishes, giving a simplified asymptotic rule: $\tau_{\text{opt}}\to a\cdot\frac{\text{Tr}({\bm{M}}_{11}^{\top}{\bm{M}}_{11})}{\text{Tr}({\bm{M}}_{11})}.$
Note that under the considered training distribution, we have $\mathrm{Tr}({\bm{M}}_{11}^{\top}{\bm{M}}_{11})/\mathrm{Tr}({\bm{M}}_{11})\approx 1$, which implies that $\tau_{\mathrm{optimal}}\to a$ as $l\to\infty$.
4.5 Possible extensions to the general softmax setting: the moment-ratio heuristic
We propose a practical heuristic (derived in Appendix J): the optimal attention temperature is roughly proportional to the ratio of the second and first moments of pre-softmax attention scores. The expression in (19) provides further theoretical justification for that heuristic.
Indeed, for some $i\neq j$,
$\tau_{\text{opt}}$ $=a\frac{\text{Tr}\left({\bm{M}}_{11}{\bm{M}}_{11}^{T}\right)}{\text{Tr}\left({\bm{M}}_{11}\right)}+\frac{1}{l}\left(\frac{\sigma^{2}}{b}+ad\right)\frac{\text{Tr}\left({\bm{M}}_{11}{\bm{M}}_{11}^{T}\right)}{\text{Tr}\left({\bm{M}}_{11}\right)},$ (20)
$=\frac{\mathbb{E}\!\left[({\bm{z}}_{i}^{\top}{\bm{M}}{\bm{z}}_{j})^{2}\right]}{\mathbb{E}\!\left[{\bm{z}}_{i}^{\top}{\bm{M}}{\bm{z}}_{i}\right]}+\frac{1}{l}\left(\frac{\sigma^{2}}{b}+ad\right)\frac{\text{Tr}\left({\bm{M}}_{11}{\bm{M}}_{11}^{T}\right)}{\text{Tr}\left({\bm{M}}_{11}\right)},$ (21)
where we used the moments calculated in Appendix J to reach the final line. Here, since $\mathrm{Tr}({\bm{M}}_{11}^{\top}{\bm{M}}_{11})/\mathrm{Tr}({\bm{M}}_{11})\approx 1$ for the considered training distribution, this gives the approximation:
$\tau_{\text{opt}}\approx\underbrace{\frac{\mathbb{E}\!\left[({\bm{z}}_{i}^{\top}{\bm{M}}{\bm{z}}_{j})^{2}\right]}{\mathbb{E}\!\left[{\bm{z}}_{i}^{\top}{\bm{M}}{\bm{z}}_{i}\right]}}_{\text{moment-ratio}}+\underbrace{\frac{1}{l}\left(\frac{\sigma^{2}}{b}+ad\right)}_{\text{correction for small $l$}}.$ (22)
This demonstrates that the moment-ratio heuristic is not an ad-hoc rule, but a theoretically grounded approximation of the exact closed-form optimal temperature.
@@ 4.5 Possible extensions to the general softmax setting: the moment-ratio heuristic
Figure 4: Effect of attention temperature on the ICL performance of Llama2-7B (Touvron et al., 2023) on the SCIQ dataset (Welbl et al., 2017). Distribution shift is induced by injecting noisy yet “relevant” labels into in-context demonstrations following Gao et al. (2024). Panel (a) fixes the noisy ratio at 0.6; panel (b) fixes the number of in-context examples at 6. Results (averaged over 12 Monte Carlo runs) include error bars showing one standard deviation. Attention temperature of all the layers is set to $\tau\sqrt{d_{k}}$ for dimension independence, where $d_{k}$ denotes the key dimension of the corresponding layer. Furthermore, the dashed black line marks the “optimal temperature” computed from the variance-to-mean ratio of pre-softmax scores, which is an insight derived from Theorem 4.3, as explained in Appendix J. Full experimental details appear in Appendix K.
5 Experimental results
In this section, we empirically validate our theory and show that optimal attention temperature consistently enhances generalization. We begin with controlled linear regression experiments using (i) the simplified Transformer model with approximate softmax attention (7) and (ii) GPT-2 (Radford et al., 2019), which combines multi-head softmax attention with MLP layers [^2] GPT-2 results are in Appendix K.. These experiments confirm that our theoretical insights transfer from simplified to expressive architectures. Finally, we evaluate Llama2-7B (Touvron et al., 2023) on SCIQ in-context learning tasks (Welbl et al., 2017), demonstrating that temperature selection is a principled and effective lever for improving robustness in large language models.
5.1 Experiments on linear regression tasks
We consider a Transformer architecture with approximate softmax attention and no MLP layers, as analyzed in our theoretical development. Figures 2 and 3 illustrate its behavior on linear regression tasks (2). Theoretical predictions closely match empirical performance across a range of conditions, confirming the robustness of our analysis. In Figure 2, we compare the ICL performance of the model with and without applying the optimal temperature. As context length $l$ increases (Figure 2a), the model’s predictions converge to those of the Bayes-optimal linear model, validating its ICL capability. Figure 2b shows that under an input covariance shift, model performance degrades—but applying the optimal temperature restores alignment with the Bayes-optimal solution. Additionally, Figure 2c shows that the influence of task distribution shift decreases as $l$ increases.
5.2 Experiments with LLMs for in-context Q&A
@@
Figure 4 shows the results. In (a) and (b), optimal attention temperature enhances the ICL performance for various context lengths and noisy ratios (analogous to Figures 3a and 3b). Furthermore, in (b), higher noise ratios push the optimal temperature upward, matching our theoretical prediction (cf. Figure 3c). Together, these experiments demonstrate that the optimal temperature is not only theoretically motivated but also an effective tool for improving ICL robustness in real-world LLMs.
@@ 6 Conclusion
This work provides a unified theoretical and empirical account of how attention temperature governs the in-context learning (ICL) performance of pretrained Transformers under distribution shift. Using a simplified yet expressive framework based on approximate softmax attention, we analytically show how shifts in input covariance and label noise degrade ICL and derive an optimal temperature that provably minimizes generalization error. Extensive experiments on synthetic regression tasks, GPT-2, and Llama-2 validate our predictions, demonstrating that temperature selection is not a mere heuristic but a principled mechanism for improving robustness. Taken together, our results advance the theoretical understanding of Transformer behavior under distribution shift and establish attention temperature as a powerful, practical lever for building more adaptive and generalizable foundation models.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
This work is supported in part by TÜBİTAK under project 124E063 within the ARDEB 1001 program, and we gratefully acknowledge their support. S.D. is further supported by an AI Fellowship from the KUIS AI Center and a PhD Scholarship (BİDEB 2211) from TÜBİTAK.
Appendix A Extended discussion of the related work
ICL by Transformers —
The ICL capability of Transformers was first brought to prominence by (Brown et al., 2020), leading to a surge of empirical and theoretical investigations. Several works have demonstrated that ICL performance improves with model scale (Wei et al., 2022; Olsson et al., 2022; Schaeffer et al., 2023), underscoring its importance in modern AI systems. To better understand this phenomenon, synthetic tasks such as linear regression have served as controlled testbeds for analyzing ICL in Transformers (Garg et al., 2022; Zhang et al., 2024; Raventós et al., 2023). A prevailing hypothesis in recent theoretical work is that Transformers implicitly learn algorithms during pretraining, which they subsequently execute during inference (Bai et al., 2023; Li et al., 2023; Akyürek et al., 2023; Ahn et al., 2023; Von Oswald et al., 2023; Mahankali et al., 2024; Fu et al., 2024; Zhang et al., 2024; Li et al., 2024; Park et al., 2024). There remains ongoing debate over the precise nature of these learned procedures. However, our work focuses on a fundamentally different question, which is how attention temperature affects the ICL performance of pretrained Transformers under distribution shifts.
Appendix B Derivation of Bayes-optimal ridge estimator for $\bm{w}$}
@@
Appendix C Derivation of approximate softmax
@@
To obtain an approximation, we expand around the origin ${\bm{z}}=\mathbf{0}$ using a first-order Taylor series:
@@
Substituting back, we obtain the approximate softmax:
@@
$=:\widehat{\text{softmax}}({\bm{z}}).$ (41)
This derivation yields the approximate softmax attention formulation in (7). From a practical standpoint, approximate softmax attention mechanisms have been empirically evaluated and shown to achieve performance comparable to standard softmax attention (Han et al., 2024).
Appendix D Temperature effects for softmax and approximate softmax
@@
In the approximate case, temperature similarly acts as an inverse scaling of the variance of the output components, capturing the limit $\tau\to\infty$ (all elements equal to $1/l$). For $\tau\to 0^{+}$, approximate softmax also reflects the maximal variance, but it does not produce a true one-hot distribution. Thus, approximate softmax closely mirrors the temperature behavior of softmax, except in the degenerate limit $\tau\to 0^{+}$, which is not of practical relevance in this work.
To further illustrate these effects, Figure 1 compares softmax and approximate softmax across different temperatures. The figure demonstrates that approximate softmax faithfully captures the variance effect of temperature: the variance of the output components is inversely proportional to $\tau$. Moreover, as $\tau\to\infty$, both softmax and approximate softmax concentrate around $1/l$, whereas linear attention with temperature scaling does not. Overall, the output distributions of softmax and approximate softmax are highly similar, except at very small values of $\tau$, where approximate softmax may yield negative components while softmax tends toward sparsity with many zeros. By contrast, linear attention with temperature scaling produces qualitatively different distributions. This comparison highlights the advantage of approximate softmax as a faithful surrogate for analyzing temperature effects relevant to softmax.
Appendix E Expanded form of approximate softmax attention
Using block matrix notation, the prediction from the approximate softmax attention model can be expanded as: